OpenStudy (anonymous):

Evaluate cot(arcsin1/x-1)

OpenStudy (anonymous):

for x=?

OpenStudy (anonymous):

I have no idea i think it just needs to be simplified..

OpenStudy (anonymous):

is this \[\cot(\arcsin(\frac{1}{x-1}))\]?

OpenStudy (anonymous):

if so we can do it so let me know

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

wow you are back?

OpenStudy (anonymous):

make a triangle. "opposite side" label 1 hypotenuse label x-1 find the adjacent side by pythagoras get \[\sqrt{(x-1)^2-1}=\sqrt{x^2-2x}\]

OpenStudy (anonymous):

cotangent is adjacent over opposite so it is just \[\sqrt{x^2-2x}\]

OpenStudy (anonymous):

I know it took me a while but omg thanks!!! yay